A man spent one fourth of his money on clothes and two fifth on shops. He had 400 left. What was his original money.
Answer by 4198181(13)

Find common denominator of 1/4 and 2/5 which is 20,
5/20+8/20=13/20
The man spent 13/20 of his money, and has 7/20 left which is $400
Let x represent the original amount:
7/20(x)=400
7(x)=8000
x=1142.857143
Answer: His original money was about $1142.86